the sum of two numbers is 39. When the smaller number is doubled and the larger is increased by 3, the resulting numbers are equal. What are the numbers?
Found 2 solutions by nerdybill, richwmiller:
Answer by nerdybill(7384)

the sum of two numbers is 39. When the smaller number is doubled and the larger is increased by 3, the resulting numbers are equal. What are the numbers?
.
Let x = smaller number
and y = larger number
.
x + y = 39
2x = y+3
.
Reordering the pair of equations above:
x + y = 39
2x - y = 3
.
Adding the two equations together:
x + y = 39
2x - y = 3
-----------
3x = 42
x = 14 (smaller number)
.
Larger number:
x + y = 39
14 + y = 39
y = 39-14
y = 25 (larger number)

Answer by richwmiller(17219)

a+b=39
2a=b+3
a=14
b=25
check
2*14=25+3
28=28
ok